# Tredoku Patterns

### v2

@@ 1 Introduction
Donald Preece asked: What arrangements of diamonds might be used in tredoku puzzles? He called these arrangements tredoku patterns. The question can be thought of as being part of the theory of diamond tilings (often called lozenge tilings) studied as part of combinatorial geometry, algebraic combinatorics, and physics. The history of such tilings in combinatorics goes back as least as far as 1896, where Percy MacMahon [4] conjectured a formula for the number of diamond tilings of a hexagonal region of the plane; he established this formula in 1916 [5]. In physics, diamond tilings are studied as simple instantiations of the 3-dimensional Ising model; in particular there have been some beautiful results on random tilings motivated by this model [2]. Diamond tilings, and the related domino tilings, are the basic examples which are generalised by the theory of dimers [3]. One interesting aspect of tredoku patterns is the introduction of the concept of a run of tiles (which we define below) to this research area.
Donald Preece gave two talks [7, 8] on this topic in 2013 in which he defined tredoku patterns, and provided many examples. (The precise definition is below. In Appendix A, the abstract from [7] is reproduced, which contains the definition in his own words.) In these talks, he provided various constructions for tredoku patterns, and proposed questions and conjectures concerning which parameters are possible. Prof Preece died in January 2014, before he was able to complete his work on tredoku patterns. This paper solves (in the affirmative) the conjectures he proposed.
I am working from the notes I took when attending his talk in July 2013 [7]. Prof.  Preece distributed copies of his notes on the topic after his talk in October [8], in order to encourage attendees to finish his work. Unfortunately, I did not take a copy myself, and I am unable to locate any remaining copies. If anyone reading this paper does have a copy of these notes, please get in touch with me: I would be very interested to see their contents.
@@
Theorem 2 is visualised by Figure 2, where we plot $\tau$ against $\ell$, and where the shaded squares are parameters where tredoku patterns exist. I believe Donald Preece had a proof of Theorem 2 modulo his conjectures above, so the material in Sections 23 and 4 below is a reconstruction of his work. His conjectures are solved in Sections 5 and 6.
@@ 2 Examples of tredoku patterns
Figure 11: Tredoku patterns with $5$ tiles and $9$ tiles and four leaves. One leaf is shaded in the $9$ tile pattern.
@@ 4 Some consequences of the connectivity conditions
Clearly $\mathcal{U}^{\prime}$ is path connected: paths from any points $x,y\in\mathcal{U}^{\prime}$ to $\mathcal{B}$ can be obtained by truncating paths to a point $\mathcal{C}$, and $\mathcal{B}$ is path connected. Let $\pi$ be a loop in $\mathcal{U}^{\prime}$. By the simplicial approximation theorem, by replacing $\pi$ by a homotopic loop if necessary, we may assume that $\pi$ is piecewise linear. We may contract $\pi$ to a point via some piecewise linear homotopy $h:[0,1]\times[0,1]\rightarrow\mathcal{U}$, since $\mathcal{U}$ is simply connected. The set $\{(x,t)\in[0,1]\times[0,1]:h(x,t)\not\in\mathcal{U}^{\prime}\}$ is a finite set of open polygons, bounded by elements in $\mathcal{B}$. Since $\mathcal{B}$ is a single edge, we may modify the homotopy on each such polygon using a piecewise-linear function in $\mathcal{B}$ to produce a homotopy $h^{\prime}:[0,1]\times[0,1]\rightarrow\mathcal{U}^{\prime}$ in $\mathcal{U}^{\prime}$ that contracts $\pi$ to a point. So $\mathcal{U}^{\prime}$ is simply connected, as required. ∎
@@ 5 Classifying tredoku patterns with $\tau=2\rho+1$
Note that many leaves $t$ will not possess tiles $s_{1}$ and $s_{2}$ satisfying the conditions above. For example, in Figure 11 only the shaded leaf possesses neighbouring tiles $s_{1}$ and $s_{2}$ of the right form. In this case, there is a unique choice for the tile northwest of the shaded leaf, but either choice for the tile southeast of the leaf will work. The $2$-leaf extension in this example is also a tredoku pattern, with $7$ tiles and $5$ leaves. As an aside, we state the following lemma, which shows that the $2$-leaf extension of a tredoku pattern is always a tredoku pattern.

### v3

@@ 1 Introduction
Donald Preece asked: What arrangements of diamonds might be used in tredoku puzzles? He called these arrangements tredoku patterns. The question can be thought of as being part of the theory of diamond tilings (often called lozenge tilings) studied as part of combinatorial geometry, algebraic combinatorics, and physics. The history of such tilings in combinatorics goes back as least as far as 1896, where Percy MacMahon [4] conjectured a formula for the number of diamond tilings of a hexagonal region of the plane; he established this formula in 1916 [5]. (MacMahon phrased the problem in terms of plane partitions: two-dimensional arrays of integers which are non-decreasing in both rows and columns. These objects can be visualised as piles of cubes, the height of each pile corresponding to the integers in the array. The appropriate projection of these cubes to a plane gives a diamond tiling.) In physics, diamond tilings are studied as simple instantiations of the 3-dimensional Ising model; in particular there have been some beautiful results on random tilings motivated by this model [2]. Diamond tilings, and the related domino tilings, are the basic examples which are generalised by the theory of dimers [3]. One interesting aspect of tredoku patterns is the introduction of the concept of a run of tiles (which we define below) to this research area.
Donald Preece gave two talks [7, 8] on this topic in 2013 in which he defined tredoku patterns, and provided many examples. (The precise definition is below. In Appendix A, the abstract from [7] is reproduced, which contains the definition in his own words.) In these talks, he provided various constructions for tredoku patterns, and proposed questions and conjectures concerning which parameters are possible. Prof.  Preece died in January 2014, before he was able to complete his work on tredoku patterns. This paper solves (in the affirmative) the conjectures he proposed.
@@
Theorem 2 is visualised by Figure 2, where we plot $\tau$ against $\ell$, and where the shaded squares are parameters where tredoku patterns exist. Donald Preece had a proof of Theorem 2 modulo his conjectures above, so the material in Sections 23 and 4 below is a reconstruction of his work [^2] After his talk in October [8], Prof.  Preece distributed copies of his notes on tredoku tilings in order to encourage attendees to finish his work. The original version of this paper was submitted with a plea for anyone with these notes to get in touch, and (remarkably!) Martin Ridout contacted me to say that he had a folder containing Prof.  Preece’s notes on tredoku patterns, including what is almost certainly his handout. (Martin was responsible for sorting through Prof.  Preece’s notes after he passed away.) My thanks to Martin for passing this folder on to me. [9]. His conjectures are solved in Sections 5 and 6. After the first version of this paper was submitted, Martin Ridout sent me details of a computer program he had written which enumerates all small tredoku patterns, which in particular shows that the first of Preece’s conjectures holds and provides an alternative proof of Theorem 18 below. He hopes to make this data, and a description of the programming techniques, available in the next few months.
@@ 2 Examples of tredoku patterns
Figure 11: Tredoku patterns with $5$ tiles and $8$ tiles and four leaves. One leaf is shaded in the $8$ tile pattern.
@@ 4 Some consequences of the connectivity conditions
Clearly $\mathcal{U}^{\prime}$ is path connected: paths from any points $x,y\in\mathcal{U}^{\prime}$ to $\Delta$ can be obtained by truncating paths to a point $\mathcal{C}$, and $\Delta$ is path connected. Let $\pi$ be a loop in $\mathcal{U}^{\prime}$. By the simplicial approximation theorem, by replacing $\pi$ by a homotopic loop if necessary, we may assume that $\pi$ is piecewise linear. We may contract $\pi$ to a point via some piecewise linear homotopy $h:[0,1]\times[0,1]\rightarrow\mathcal{U}$, since $\mathcal{U}$ is simply connected. The set $\{(x,t)\in[0,1]\times[0,1]:h(x,t)\not\in\mathcal{U}^{\prime}\}$ is a finite set of open polygons, bounded by elements in $\Delta$. Since $\Delta$ is a single edge, we may modify the homotopy on each such polygon using a piecewise-linear function in $\Delta$ to produce a homotopy $h^{\prime}:[0,1]\times[0,1]\rightarrow\mathcal{U}^{\prime}$ in $\mathcal{U}^{\prime}$ that contracts $\pi$ to a point. So $\mathcal{U}^{\prime}$ is simply connected, as required. ∎
@@ 5 Classifying tredoku patterns with $\tau=2\rho+1$
Note that many leaves $t$ will not possess tiles $s_{1}$ and $s_{2}$ satisfying the conditions above. For example, in Figure 11 only the shaded leaf possesses neighbouring tiles $s_{1}$ and $s_{2}$ of the right form. In this case, there is a unique choice for the tile northwest of the shaded leaf, but either choice for the tile southeast of the leaf will work. The $2$-leaf extension in this example is also a tredoku pattern, with $10$ tiles and $5$ leaves. As an aside, we state the following lemma, which shows that the $2$-leaf extension of a tredoku pattern is always a tredoku pattern.